OpenStudy (anonymous):

14+(-6)+(-27)=? A.47 B.19 C.-19 D.-47

OpenStudy (anonymous):

@inowalst

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@Rebekah2001

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

ill help u

OpenStudy (anonymous):

The answer is C. -19.

OpenStudy (sleepyjess):

Stew, please do not give direct answers

OpenStudy (anonymous):

could someone from here please help me if you don't mind.

OpenStudy (anonymous):

14+(-6)+(-27)=? = 14-6-27 starting from left to right =8-27 =-19

OpenStudy (anonymous):

c